# Improving Lesion Segmentation in Medical Images by Global and Regional Feature Compensation

### v1

@@ 4.4.2 Ablation Study
Adding the GCU to skip connections improved DSC and Precision by 1.3% and 1.7%, respectively from the UNet++ baseline (Table 2). The RCU introduced two key operations: (1) applying cross-attention between decoder feature maps and the MRM, and (2) generating patch-based importance scores. Building on the GCU results, adding cross-attention with MRM further improved DSC and Precision by another 1.5% and 1.3%. Incorporating patch-based importance scores provided an additional boost of 1.4% and 1.3%.
Overall, GRCSF improved DSC, IoU, F1, Precision, and Recall by 4.1%, 3.3%, 1.8%, 2.1%, and 4.2%, from the baseline, respectively. These results demonstrate that the integration of all proposed modules significantly enhances the framework’s segmentation performance, validating the design choices of GRCSF.
@@ 5 Discussion
CNN-based models like U-Net and UNet++ rely on fixed skip connections that cannot adaptively refine feature maps, resulting in the loss of detailed features. Their fixed receptive fields further limit their ability to capture global contextual information required for segmenting lesions of varying scales and locations. Consequently, U-Net tends to over-segment lesions or miss small lesions altogether. While UNet++ mitigates over-segmentation compared to U-Net, it still struggles to detect challenging lesions with sufficient sensitivity. Similarly, SPiN improves resolution by upsampling input images but struggles with accurate segmentation due to its tendency to over-segment, leading to false positives. SPiN lacks mechanisms to dynamically focus on critical regions, which further limits its precision in complex cases. GRCSF effectively addresses the limitations of the above-mentioned methods in challenging lesion segmentation tasks. It significantly improves sensitivity to small and low-contrast lesions by feature compensation. GRCSF also resolves the blurred edges and ambiguous segmentation common in the output produced by other methods. Additionally, by dynamically focusing on lesion-specific regions, GRCSF minimizes false positives, enhancing the reliability of segmentation results in complex cases. Transformer-based models such as TransUNet and SwinUNet attempt to capture long-range dependencies and enhance global spatial information. However, their patch-based designs often struggle to capture detailed context around small lesions and fail to provide sufficient edge precision, resulting in high false-negative rates. Additionally, these models are prone to overfitting and suboptimal results on small datasets such as orCaScore, reducing their practicality and generalizability. In contrast, GRCSF integrates global and regional features effectively without requiring complex parameter tuning. It achieves higher precision in lesion segmentation while avoiding over-segmentation. The model’s superior performance, even without post-processing of orCaScore results, further highlights its practicality for real-world applications where post-processing conditions are often limited.
@@
In particular, the ATLAS dataset posed unique challenges due to significant image variability, including differences in imaging protocols, scanner types, and cohort demographics across medical centers. In our study, certain subsets involved training on a few cohorts while testing on entirely different cohorts, reflecting differences in cohort characteristics. Despite these challenges, GRCSF demonstrated exceptional robustness, consistently achieving higher segmentation performance across all four subsets, thereby validating tis strong generalizability under diverse training and testing conditions.
@@
The observed improvements in the ablation study highlight the complementary roles of global and regional compensation in GRCSF. The effective utilization of pixel-level feature similarity from different sources helps the model focus on regions that are often overlooked. The 50% and 75% mask ratios in MAE generate complementary MRM, which are adaptively integrated through learnable weights. This integration dynamically highlights informative regions while suppressing irrelevant ones, thereby mitigating false negatives and false positives caused by reconstruction errors. By leveraging these MRM through the RCU, GRCSF prioritizes regions with high lesion likelihood and compensates for missed feature representations, which is particularly effective for low-contrast and small lesions. This demonstrates the added value of MRM in comparison to baseline networks, which exhibited suboptimal performance in these challenging scenarios.
@@
Unlike traditional SSL methods that use MAE for pre-training, our method directly utilizes the reconstructed images to generate SSL residual maps. While involving domain and high-level image features, preserving more low-level and pixel-level information. This method aligns with the learnable and consistent anatomical structures across patients. These benefits are not achievable through pre-trained encoders alone. Moreover, averaging five runs of MRM generation mitigates the variability introduced by random masking, ensuring reliable guidance for the backbone. This strategy significantly improves segmentation accuracy and robustness, offering a novel way to integrate SSL into medical image analysis.
@@ 6 Conclusion
In this work, we propose GRCSF, a dual-feature compensation framework that leverages a U-shaped backbone and SSL residual maps to improve lesion segmentation. We evaluate its effectiveness using two example datasets representing low-contrast and small lesion segmentation tasks. The GCU mitigates information loss caused by downsampling in the encoder, enhancing global feature representation to improve boundary delineation and small lesion detection. The RCU integrates complementary pixel-level features from MRM, employing a patch-based importance scoring mechanism to localize lesions more effectively while reducing false positives. Together, GRCSF provides a robust and efficient solution, demonstrating improvement in challenging medical imaging segmentation tasks.
Limitation: Our method relies on MAE-generated residual maps, which can be computationally demanding. In addition, though the method performs well on two datasets, variations in imaging protocols across medical centers still present challenges for broader application. Future efforts should focus on reducing model complexity and improving generalizability by introducing model compression and domain adaptation techniques.

### v2

@@ 4.4.2 Ablation Study
Adding the GCU to skip connections improved DSC and Recall by 1.3% and 1.7%, respectively from the UNet++ baseline (Table 3). The RCU introduced two key operations: (1) applying cross-attention between decoder feature maps and the MRM, and (2) generating patch-based importance scores. Building on the GCU results, adding cross-attention with MRM further improved DSC and Recall by another 1.4% and 1.3%. Incorporating patch-based importance scores provided an additional boost of 1.4% and 1.2%.
Overall, GRCSF improved DSC, IoU, Precision, and Recall by 4.1%, 3.3%, 2.1%, and 4.2%, from the baseline, respectively. These results demonstrate that the integration of all proposed modules significantly enhances the framework’s segmentation performance, validating the design choices of GRCSF.
Module Method Dice IoU Precision Recall Model Parameters
Residual Map in RCU Grad-CAM 0.476 0.359 0.470 0.566 42.85 M
SimMIM Residual Map (AbsDiff) 0.553 0.421 0.537 0.663 42.85 M
MAE Residual Map (MSE) 0.547 0.422 0.541 0.648 42.85 M
MAE Residual Map (SSIM) 0.529 0.404 0.526 0.633 42.85 M
MAE Residual Map (AbsDiff) (Ours) 0.581 0.448 0.595 0.645 42.85 M
GCU UNet++ with PAM 0.490 0.399 0.558 0.613 72.21M
UNet++ with AG skip connections 0.521 0.413 0.513 0.618 40.12M
SAM in GCU 0.526 0.415 0.548 0.635 42.20M
GCU (Ours) 0.534 0.408 0.560 0.615 42.27M
Segmentation Backbone UNet with pre-trained MAE Encoder 0.517 0.395 0.559 0.583 304.93M
GRCSF with MobileNet Encoder 0.472 0.358 0.490 0.549 35.13M
GRCSF (Ours) 0.581 0.448 0.595 0.645 42.85 M
Table 7: Impact of different module design choices on segmentation performance, evaluated on one ATLAS test subset. The best results within each module group are highlighted in bold.
4.4.3 Design Choices of GRCSF Modules
The results in Table 4 indicate that our framework achieves the best performance when residual maps are generated using five MAE iterations. Although this configuration introduces an additional 17.53 seconds compared to using a single iteration, it yields improvements of 6.1% in Dice and 4.8% in IoU. Table 5 shows that, when using a single mask ratio, 50% and 75% are the most effective. Combining these two ratios leads to the best overall performance in terms of Dice and IoU. As shown in Table 6, the cross-attention configuration with patch sizes of 8×8, 8×8, and 16×16 outperforms other combinations.
Figure 11: Visual comparison of residual maps produced by different methods: MAE residuals computed with absolute difference (AbsDiff), MSE, and structural similarity (SSIM); SimMIM residuals at 50% and 75% mask ratios using AbsDiff; and the Grad-CAM map from UNet. Please zoom in for a clearer view.
The residual maps generated by SimMIM and MAE using AbsDiff, MSE, SSIM, and Grad-CAM are visualized in Fig. 11. Among them, the MAE residual map obtained with AbsDiff achieves the highest Dice score, followed by SimMIM. Replacing the GCU with PAM and AG skip connections does not improve performance. Similarly, substituting the SE module in our GCU design with SAM does not yield better results. Moreover, replacing the encoder of UNet++ with a lightweight architecture such as MobileNet leads to a significant drop in performance. Additionally, implementing MAE pre-trained encoder for segmentation does not demonstrate notable improvements. The results are shown in Table 7.
@@ 5 Discussion
CNN-based models like U-Net and UNet++ rely on fixed skip connections that cannot adaptively refine feature maps, resulting in the loss of detailed features. Their fixed receptive fields further limit their ability to capture global contextual information required for segmenting lesions of varying scales and locations. Consequently, U-Net tends to over-segment lesions or miss small lesions altogether. While UNet++ mitigates over-segmentation compared to U-Net, it still struggles to detect challenging lesions with sufficient sensitivity. Similarly, SPiN improves resolution by upsampling input images but struggles with accurate segmentation due to its tendency to over-segment, leading to false positives. SPiN lacks mechanisms to dynamically focus on critical regions, which further limits its precision in complex cases. GRCSF effectively addresses the limitations of the above-mentioned methods in challenging lesion segmentation tasks. It significantly improves sensitivity to small and low-contrast lesions by feature compensation. GRCSF also resolves the blurred edges and ambiguous segmentation common in the output produced by other methods. Additionally, by dynamically focusing on lesion-specific regions, GRCSF minimizes false positives, enhancing the reliability of segmentation results in complex cases. Transformer-based models such as TransUNet, TransFuse, SwinUNet and UTNet attempt to capture long-range dependencies and enhance global spatial information. However, their patch-based designs often struggle to capture detailed context around small lesions and fail to provide sufficient edge sensitivity, resulting in high false-negative rates. Additionally, due to their high model complexity, these methods are prone to overfitting and yield suboptimal results on small datasets such as orCaScore, thereby limiting their practicality and generalizability. In contrast, GRCSF integrates global and regional features effectively without requiring complex parameter tuning. It achieves higher recall in lesion segmentation while avoiding over-segmentation. The model’s superior performance, even without post-processing of orCaScore results, further highlights its practicality for real-world applications where post-processing conditions are often limited. Notably, SAM does not outperform our method on any of the three datasets. However, it demonstrates effectiveness in refining high-contrast and irregular lesions, such as lung tumors, as evidenced by its improved performance over UNet++ on the MSD Lung Tumor dataset when refining the bounding boxes derived from UNet++ predictions. In contrast, it struggles with low-contrast brain lesions, where the same prompts result in worse performance than UNet++. Moreover, SAM lacks the ability to correct false positives at the lesion level; as a result, when prompts frequently include such cases, it leads to lower precision and higher FPR and VOSE. Its performance could potentially be improved through few-shot learning and advanced prompt engineering strategies.
@@
In particular, the ATLAS dataset posed unique challenges due to significant image variability, including differences in imaging protocols, scanner types, and cohort demographics across medical centers. In our study, certain subsets involved training on a few cohorts while testing on entirely different cohorts, reflecting differences in cohort characteristics. Despite these challenges, GRCSF demonstrated exceptional robustness, consistently achieving higher segmentation performance across all four subsets, thereby validating its strong generalizability under diverse training and testing conditions. While our method incorporates residual maps generated by MAE, which introduces additional computational cost, this increase remains within practical limits. Specifically, GRCSF requires only 23.51 and 35.97 seconds more than the fastest baseline UNet per patient on the ATLAS and MSD Lung Tumor datasets, respectively, yet achieves Dice improvements of 5% and 6%. In time-sensitive scenarios, the full pipeline can process a patient in approximately 1 minute on a single A6000 GPU, offering a favorable trade-off between accuracy and efficiency for routine deployment.
@@
The observed improvements in the ablation study highlight the complementary roles of global and regional compensation in GRCSF. The effective utilization of pixel-level feature similarity from different sources helps the model focus on regions that are often overlooked. The 50% and 75% mask ratios in MAE generate complementary MRM, which are adaptively integrated through learnable weights and cross-attention mechanism. This integration dynamically highlights informative regions while suppressing irrelevant ones, thereby mitigating false negatives and false positives caused by reconstruction errors. By leveraging these MRM through the RCU, GRCSF prioritizes regions with high lesion likelihood and compensates for missed feature representations, which is particularly effective for low-contrast and small lesions. This demonstrates the added value of MRM in comparison to baseline networks, which exhibited suboptimal performance in these challenging scenarios.
@@
Unlike traditional SSL methods that use MAE for pre-training, our method directly utilizes the reconstructed images to generate SSL residual maps. While involving domain and high-level image features, preserving more low-level and pixel-level information. This method aligns with the learnable and consistent anatomical structures across patients. These benefits are not achievable through pre-trained encoders alone. Although the MAE pre-training configuration outperforms the original UNet baseline, it still falls short of the performance achieved by the full GRCSF pipeline. Moreover, using a fixed MAE encoder limits architectural flexibility. In contrast, components such as the GCU can be integrated into U-shaped backbones but become more complex to implement with a frozen MAE encoder. Moreover, averaging five runs of MRM generation mitigates the variability introduced by random masking, ensuring reliable guidance for the backbone. This strategy significantly improves segmentation accuracy and robustness, offering a novel way to integrate SSL into medical image analysis. An alternative of MAE MRM is SimMIM MRM, which appears to produce smoother reconstructions and visually sharper structural details, which makes the resulting residual map less suitable for lesion segmentation, especially for small lesions. This is because the SimMIM reconstruction allows normal anatomical structures to dominate the residual, obscuring the regions that contain small lesions. In contrast, MAE captures global morphological features throughout the dataset and leaves larger reconstruction errors over abnormal tissues, which more effectively highlight lesion areas.
@@ 6 Conclusion
In this work, we propose GRCSF, a dual-feature compensation framework that leverages a U-shaped backbone and SSL residual maps to improve lesion segmentation. We evaluate its effectiveness using three example datasets representing low-contrast and small lesion segmentation tasks. The GCU mitigates information loss caused by downsampling in the encoder, enhancing global feature representation to improve boundary delineation and small lesion detection. The RCU integrates complementary pixel-level features from MRM, employing a patch-based importance scoring mechanism to localize lesions more effectively while reducing false positives. Together, GRCSF provides a robust and efficient solution, demonstrating improvement in challenging medical imaging segmentation tasks.
Limitation: The overall computational cost, including residual map generation and the use of a moderately sized backbone, may present challenges for large-scale deployment or routine clinical use. In our future work, we will aim to reduce model complexity by exploring efficient strategies for residual map generation, for example, replacing random masking with a more targeted approach that applies masks to regions of interest. In addition, the current UNet++ backbone could be substituted with a lightweight yet high-performance alternative to further reduce training and inference cost.
7 Acknowledgments
This work was supported by the Australian Research Council Discovery project (DP200103748), NHMRC Investigator (APP2017023) and in part by Startup funds from The University of Newcastle.